Health Insurance. Determine the subsidy. ($millions)
D: P = 5,000 - Q
S: P = 400 + Q
Co-pay: P = 60
1
Expert's answer
2016-05-24T12:42:02-0400
D: P = 5,000 - Q
S: P = 400 + Q
In equilibrium Qd = Qs and Pd = Ps, so:
5,000 - Q = 400 + Q,
2Q = 4,600,
Q = 2,300 units.
P = 5,000 - 2,300 = $2,700.
If co-pay is P = 60, then the amount of the subsidy is 2,300*60 = $13,800.